a firm sells a 20 year bond for a premium of 25,000 over its 250,000 face value. bond's coupon rate is 7% and they use straight line amortization. what is their interest expense on the bond each year.
Answer #1

Annual interest payment = Par value of bonds x Interest rate

= 250,000 x 7%

= $17,500

Premium on bonds payable =$25,000

Bond life = 20 years

Annual amortization of bond premium = Premium on bonds payable/Bond life

= 25,000/20

= $1,250

Annual interest expense =Annual interest payment- Annual amortization of bond premium

= 17,500-1,250

= $16,250

Interest expense on the bond each year = $16,250

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!
Answer #2

SOLUTION :


Face value of the bond = 250000 


Premium = 25000


Maturity period of the bond = 20 years.


Coupon rate is 7% annually.


Hence, coupon amount per year 

= Interest expense on the bond

= 250000 * 0.07

= 17500 


Premium of 25000 is to be amortised for 20 years


As per straight line method amortisation per year

= 25000 / 20

= 1250


So, total expense (considered as interest) per year

= Coupon amount - amortised premium

= 17500 -  1250

= 16250 (ANSWER).

answered by: Tulsiram Garg